OpenStudy (anonymous):

X x X -4X -13=0 HELP PLEASE

OpenStudy (anonymous):

Well: x*x-4x-13=0 multiply the x's x^2-4x-13=0 now we need to factor, it doesn't look like there will be a simple answer so we will need to use the quadratic formula: x=-b\[x=(-b \pm \sqrt{b^2 - 4ac})/2a\] in your case: a=1, b=-4, and c=-13 so x= \[(4 \pm \sqrt{16-4*1*(-13)})/2*1\] now we'll just continue to simplify x= \[(4 \pm \sqrt{16+52})/2\] x= \[(4 \pm \sqrt{68})/2\] x= \[(4 \pm \sqrt{17*4})/2\] x=\[(4 \pm 2\sqrt{17})/2\] x=\[(2 \pm \sqrt{17})\] and that is your final answer: x= \[(2 + \sqrt{17})\], \[(2 - \sqrt{17})\] let me know if you have more questions